# Working Backwards: Learning to Place by Picking

### v2

@@ Abstract
We present placing via picking (PvP), a method to autonomously collect real-world demonstrations for a family of placing tasks in which objects must be manipulated to specific contact-constrained locations. With PvP, we approach the collection of robotic object placement demonstrations by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems. Specifically, we obtain placing demonstrations from a set of grasp sequences of objects initially located at their target placement locations. Our system can collect hundreds of demonstrations in contact-constrained environments without human intervention by combining two modules: tactile regrasping and compliant control for grasps. We train a policy directly from visual observations through behavioral cloning, using the autonomously-collected demonstrations. By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table). We validate our approach in home robotic scenarios that include dishwasher loading and table setting. Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
@@ I Introduction
Imitation learning (IL) provides an appealing, simple, and practical option by learning end-to-end control policies from expert demonstrations [1, 2]. The general approach of training expressive models on large and diverse datasets has proven to be quite effective outside of robotics. While there have been many attempts to transfer this paradigm to robotic systems [3, 4], robots have yet to achieve similar successes. A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect the large amount of expert demonstrations in the physical world.
Having said this, many recent methods attempt to streamline [5, 6] or automate [7, 8, 9, 10] data collection. As highlighted by the authors of [9], the lack of robust and diverse autonomous data collection policies is a major limitation to scaling up IL. In this work, we are interested in investigating autonomous data collection for robotic object placement. As opposed to prior research, we tackle the full placing problem, without restricting our task to simple, small objects placed on flat, horizontal surfaces with no environment contact constraints. We assert that object and environment contact constraints can be a valuable, natural guide for learning how to place if handled properly.
We propose a novel approach named placing via picking (PvP) that automates the collection of expert demonstrations for a large subset of contact-constrained placing problems. We do so by taking advantage of a powerful grasp planner [11], tactile sensing, and compliant control. PvP is a self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems. Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory. We generate demonstration data for object placement in a self-supervised manner—the picking phase provides the training supervision for the placing task. While appearing deceptively easy at first, we highlight the crucial importance of two modules, compliant control for grasping (CCG) and tactile regrasping (TR). CCG and TR enables (1) robust and uninterrupted pick-and-place and (2) leveraging environment contact-constraints as guides to inform object placement. LABEL:fig:title provides a visual summary of our approach. Our main contributions are:
@@
    a compliant grasping and a tactile-driven regrasping strategy to achieve uninterrupted pick and place data collection without human intervention in contact-constrained environments;
@@ II Related Work
In this section, we review previous approaches to automate robot data collection in the real world and other work that generally leverage the idea of “working backwards”.
@@ II-A Automatic Data Collection
A large number of existing policy learning approaches use prior policies to automate the data collection process. Examples include using trajectory optimization [12, 13] and, closer to our work, simple scripted pick-and-place policies [7, 8, 9] as forms of supervision. The authors of [7] use a scripted pick-and-place policy to bootstrap the initial dataset used for off-policy deep reinforcement learning with a real-world manipulator. In [9], the authors present AutoRT, a method that uses a combination of scripted, learned and tele-operated data collection policies to efficiently collect a large amount of language-specified manipulation demonstration data. Using a combination of a large language model (LLM) and visual language models (VLMs), AutoRT observes scenes in the wild and comes up with plausible tasks. Once a task is defined and chosen, AutoRT samples a data collection policy to attempt the task and collect real world data. In our work, we focus on developing a method for collecting demonstrations of placement with larger and more complex objects (i.e., objects that require a planner to grasp) in environments with potential contact constraints. Unlike prior work, we do not restrict the placing task to simpler small objects on flat horizontal surfaces. The authors of [9] mention the lack of robust and diverse real-world autonomous collection policies as a limitation of AutoRT; Our work is complementary with many of the prior frameworks and PvP could be used as a prior policy within them.
@@ II-B Working Backwards
The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [14] and in learning visual predictive models for control [15, 16, 17]. Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [18, 19, 20]. These methods consider objects that are assumed to be in their desired place configuration during data collection. The manipulator can then pick an object and move back to the initial goal location to generate data. However, we focus on placement as opposed to insertion as our task. Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple, generally larger objects and place poses involved in the task. Furthermore, unlike prior work [18, 19, 20], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning. Our approach also resembles [21], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner. However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement. In contrast, the self-supervised data collection procedure introduced in [21] uses a 2D open loop pick-and-place action primitive.
@@ II-C Robotic Manipulator Placing
Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [22, 23, 24], while placing is relatively less explored.
Previous approaches have explored the use of motion planning combined with place-specific objectives [25, 26]. For example, in [25], the authors demonstrate a hierarchical sampling-based motion planner that first finds poses which satisfy constraints (e.g., reachability) and then performs a local optimization for user-specified objectives (e.g., clearance). The place problem has also been studied from the perspective of geometric modelling [27], where the planar surfaces of a discretized object model are matched to planar patches in the environment. Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time.
Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [28] and imitation learning (IL) [29]. In [29], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing. The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in tandem with any methods that require expert demonstrations.
@@ III Placing via Picking
We first present an overview of our self-supervised data collection method, PvP, in Section III-A and then describe a method for noise augmentation during data collection in Section III-D. Next, we provide details about the procedure used to train our IL policy in Section III-E.
@@ III-A Self-supervised Data Collection
[…]
Figure 1: The four steps involved in PvP, our autonomous demonstration data collection process for placing. We (a) generate grasps with an off-the-shelf grasp planner [11], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
PvP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving, and (4) placing by reversing; these steps are visualized in Figure 1. During training, the objects of interest are initially in their goal positions.
@@ III-A1 Grasp Planning
[…]
Figure 2: The steps involved in our language-driven grasp planning pipeline. We use (a) Grounding-Dino [30] for object bounding box detection based on text and (b) Segment Anything [31] on the cropped images to produce object specific masks. We then use (c) Contact-graspnet [11] for grasp generation on only the segmented objects (i.e., masked areas). Top: using “green plate” and “blue plate” as the targets for data collection. Bottom: using “bowl” and “mug” as the targets for data collection.
At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1. We use the grasp planner Contact-GraspNet [11] to generate $L$ candidate grasp poses $\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$. Given a set of text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”), potentially from a user or high-level LLM planner, we use Grounding-Dino [30] to find the object bounding boxes. Finally, we run Segment Anything [31] on each of the previously generated image crops from the bounding boxes to generate per-object masks. Contact-GraspNet can then use these masks to filter grasps that are not on objects of interest and keep only a pruned set of grasps $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$. We visualize an example of this process in Figure 2 for two different scenes. After pruning, we randomly select a single grasp pose, $\mathbf{T}_{\text{grasp}}$, from $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$. To approach the object of interest without collisions, we include a pregrasp pose $\mathbf{T}_{\text{pregrasp}}$ that is defined with a translational offset above the grasp pose.
@@ III-A2 Grasping
Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory. An example of a grasp sequence is shown in Figure 1. During grasping, two modules play a critical role in the robustness of PvP: compliant control for grasping (CCG) and tactile regrasping (TR).
III-B Compliant Control for Grasping
@@
We use a cartesian impedance controller for manipulation. This choice allows us to set the level of compliance of the manipulator at various stages as needed. Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values. The low compliance of the robot arm then allows the manipulator and gripper to comply with the contact constraints and minimize its applied forces against the environment. Larger contact-constrained objects generate unnecessary reaction forces when gripped without compliance. These large forces trigger emergency stops and can damage the environment, and the robot. Specific to PvP, large reaction forces also disrupt and move the object relative to the gripper ultimately leading to inaccurate placements. On the other hand, when grasping with compliance, the environment contact constraints guide the manipulator and gripper towards a natural object placement pose.
Figure 3: Tactile images from before and after a tactile regrasp (TR). Left: the plate’s contact surface area fills half of the tactile image indicating a shallow grasp. Right: after a regrasp, the contact surface area has increased indicating a stable grasp.
[…]
Figure 4: Visualization of contact surfaces in the tactile image from grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
III-C Tactile Regrasping
@@
Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip significantly once grasped. In the real world, the accumulation of errors from various sources (e.g., camera calibration, controller, learned grasp planner, and noise from the RGB-D sensor and manipulator encoders) results in grasps that are unstable (e.g., a shallow grasp) and not ideal for PvP. To mitigate this issue, we use tactile sensing (specifically, a visuotactile Finger-STS [32] sensor) to preemptively detect if a grasp is stable and perform a regrasp if needed. The Finger-STS is capable of providing multimodal feedback (i.e., both tactile and visual observations). To this end, we design our regrasping module to detect the contact surface area and calculate the relative change in EE pose required to recover a larger contact surface and, thereby, a more stable grasp. We detect regions of contact based on marker displacements and the RGB image from the Finger-STS. If an impartial contact is detected, we calculate the difference between the ideal and detected contact region and use this to calculate the commanded change in EE pose. We note that a simpler and more general random regrasp strategy could also be used. An example of this is shown in Figure 3. We visualize contact surfaces of various household items in Figure 4.
@@
In Section IV-A, we experimentally investigate the effects of including these modules through ablation studies.
III-C1 Retrieving
@@
We begin the retrieval phase when a valid, stable grasp is detected. An image of the retrieval procedure is shown in Figure 1. We store the poses of the end effector (EE) as well as the respective timestamps throughout the retrieval process. For our work, we measure joint encoder readings at a rate of 120 $\mathrm{H}\mathrm{z}$. The manipulator is set to be compliant along the rotational axes of the EE only, when commanded to move to the pregrasp pose. This allows the manipulator to better comply with any unexpected environment interactions. Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{\text{clearance}}$ centered around a set fixed pose above the scene. That is, given a set fixed pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$ to recover a clearance pose. We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection.
@@
At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose:
@@
where $\mathbf{T}_{0}=\mathbf{T}_{\text{grasp}}$, $\mathbf{T}_{M}=\mathbf{T}_{\text{clearance}}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0. To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a sparser trajectory $\bar{\tau}^{\text{expert}}_{r}$. We extract a total of $\bar{M}+1$ states where $\bar{M}=\lceil\frac{t_{M}-t_{0}}{\Delta t}\rceil$. We do this by finding the nearest pose in terms of time from the dense trajectory $\tau^{\text{expert}}_{r}$ at every $\Delta t$ interval:
@@
In this work, our visual policies operate at 5 $\mathrm{H}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
III-C2 Placing by reversing
We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory

### v4

@@ Abstract
We present placing via picking (PvP), a method to autonomously collect real-world demonstrations for a family of placing tasks in which objects must be manipulated to specific, contact-constrained locations. With PvP, we approach the collection of robotic object placement demonstrations by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems. Specifically, we obtain placing demonstrations from a set of grasp sequences of objects initially located at their target placement locations. Our system can collect hundreds of demonstrations in contact-constrained environments without human intervention using two modules: compliant control for grasping and tactile regrasping. We train a policy directly from visual observations through behavioural cloning, using the autonomously-collected demonstrations. By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table). We validate our approach in home robot scenarios that include dishwasher loading and table setting. Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of success rate and data efficiency, while requiring no human supervision.
@@ I Introduction
Imitation learning (IL) provides a scaleable, simple, and practical option to learn control policies from expert demonstrations [1, 2]. The general approach of training expressive models on large and diverse datasets has proven to be quite effective outside of robotics [3, 4]. There have been many attempts to transfer this paradigm to IL for robotic systems [5, 6], but similar successes have yet to be achieved. A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect a large number of expert demonstrations in the physical world. Having said this, many recent methods attempt to streamline [7, 8] or automate [9, 10, 11, 12] data collection. As highlighted by the authors of [11], the lack of robust and diverse autonomous data collection policies is a major limitation to scaling up IL. In this work, we are interested in investigating autonomous data collection for robotic object placement. As opposed to much prior research, we tackle the full placing problem, without restricting our task to simple, small objects placed on flat, horizontal surfaces with no environment contact constraints. We assert that object and environment contact constraints can be a valuable, natural guide for learning how to place, if handled properly.
We propose a novel approach named placing via picking (PvP) that automates the collection of expert demonstrations for a large subset of contact-constrained placing problems. We do so by taking advantage of a powerful grasp planner [13], tactile sensing, and compliant control. PvP is a self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place tasks. Given an environment with objects initially at their target locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory. We generate demonstration data for object placement in a self-supervised manner—the picking phase provides the training supervision for the placing task. While appearing deceptively easy at first, we highlight the crucial importance of two modules, compliant control for grasping (CCG) and tactile regrasping (TR). CCG and TR enable (1) robust and uninterrupted pick-and-place and (2) the use of environment contact-constraints as guides to inform object placement. LABEL:fig:title provides a visual summary of our approach. Our main contributions are
@@
    a compliant grasping and a tactile-driven regrasping strategy to achieve uninterrupted pick and place data collection in contact-constrained environments;
@@ II Related Work
In this section, we review prior work on automated robot data collection, discuss the general idea of “working backwards,” and investigate existing robotic placement strategies.
@@ II-A Automatic Data Collection
A large number of existing policy learning approaches use prior policies to automate the data collection process. Examples include using trajectory optimization [14, 15] and, closer to our work, simple scripted pick-and-place policies [9, 10, 11] as forms of supervision. The authors of [9] use a scripted pick-and-place policy to bootstrap the collection of the initial dataset used for off-policy deep reinforcement learning with a real-world manipulator. In [11], the authors present AutoRT, a method that uses a combination of scripted, learned, and tele-operated data collection policies to efficiently collect a large amount of language-specified manipulation demonstration data. Using a combination of a large language model (LLM) and visual language models (VLMs), AutoRT observes scenes in the wild and comes up with plausible tasks. Once a task is defined and chosen, AutoRT samples a data collection policy to attempt to gather real world data for the task. In our work, we focus on developing a method for collecting demonstrations of placement with larger and more complex objects (i.e., objects that require a planner to grasp) in environments with contact constraints. Unlike prior work, we do not restrict the placing task to simple small objects on flat horizontal surfaces. The authors of [11] mention the lack of robust and diverse real-world autonomous collection policies as a limitation of AutoRT; our work is complementary to many of the previous frameworks, where PvP could be used as a prior policy.
Other works have explored automatic data generation by bootstrapping from a single human demonstration [16, 17]. In [16], a robot manipulation task is modelled as having two phases: a coarse approach trajectory phase towards a bottleneck pose, followed by a fine interaction trajectory phase. The single human demonstration provides the relative pose of the manipulator with respect to the task-relevant object, from which a self-supervised data collection procedure is formulated to train a bottleneck pose predictor. In our work, we also automatically collect and augment approach trajectory data. However, we do not require any human demonstrations since we focus on placing tasks only. By doing so, we can leverage the inherent symmetry of picking and placing for self-supervised data collection through the use of a combination of tactile sensing, compliant control, and an off-the-shelf grasp planner.
@@ II-B Working Backwards
The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [18], learning visual predictive models for control [19, 20, 21], and reinforcement learning [22]. Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [23, 24, 25]. These methods consider objects that are assumed to be in their desired place configuration during data collection. The manipulator can then pick an object and move back to the initial goal location to generate training data. However, we focus on placement as opposed to insertion. Insertion policies generally only reason about local interactions. On the other hand, placement involves reasoning over a larger scene with the potential for multiple place poses and more complex objects that require planners to grasp successfully. Furthermore, unlike prior work [23, 24, 25], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning. Our approach also resembles [26], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner. However, we focus on generating demonstrations for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement. In contrast, the self-supervised data collection procedure introduced in [26] uses a 2D open loop pick-and-place action primitive.
@@ II-C Robotic Manipulator Placing
Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [27, 28, 29], while placing has been relatively less studied.
Previous approaches have explored the use of motion planning combined with place-specific objectives [30, 31]. For example, in [30], the authors demonstrate a hierarchical sampling-based motion planner that first finds poses which satisfy constraints (e.g., reachability) and then performs a local optimization for user-specified objectives (e.g., clearance). The place problem has also been studied from the perspective of geometric modelling [32], where the planar surfaces of a discretized object model are matched to planar patches in the environment. Importantly, 3D models of the objects being placed as well as the environment are assumed to be available ahead of time.
Closer to our approach, other works have trained end-to-end place policies using reinforcement learning [33] and imitation learning [34]. In [34], the authors demonstrate an inverse reinforcement learning technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing. The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in tandem with methods that require expert demonstrations.
@@ III Placing via Picking
We first present an overview of our self-supervised data collection method, PvP, in Section III-A, and then describe a method for noise augmentation during data collection in Section III-B. Finally, we provide details about the procedure used to train our IL policy in Section III-C.
@@ III-A Self-Supervised Data Collection
(a) Grasp planning.
(b) Grasping.
(c) Retrieving.
(d) Placing by reversing.
Figure 1: The four steps involved in PvP, our autonomous demonstration data collection process for placing. We (a) generate grasps with an off-the-shelf grasp planner [13]; (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp via tactile sensing; (c) retrieve the object with rotational compliance while storing the trajectory; and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
PvP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving, and (4) placing by reversing; these steps are visualized in Fig. 1. During training, the objects of interest are initially in their goal positions.
@@ III-A1 Grasp Planning
(a) Raw image.
(b) Text-based object detection.
(c) Segmentation.
(d) Pruned grasps.
Figure 2: The steps involved in our language-driven grasp planning pipeline. We use (b) Grounding-Dino [35] for object bounding box detection based on text descriptions and (c) Segment Anything [36] on the cropped images to produce object specific masks. We then use (d) Contact-GraspNet [13] for grasp generation on only the segmented objects (i.e., masked areas). Top: using “green plate” and “blue plate” as the targets for data collection. Bottom: using “bowl” and “mug” as the targets for data collection.
At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Fig. 1(a). We use the grasp planner Contact-GraspNet [13] to generate $L$ candidate grasp poses $\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in\mathrm{SE}(3)$ for all $i\in\{0,...,L\}$. Given a set of text-based descriptions of the objects of interest (e.g., “plates,” “green plate,” “cup” or, more generally, “objects”), potentially from a user or high-level LLM planner, we use Grounding-Dino [35] to find the object bounding boxes. Finally, we run Segment Anything [36] on each of the previously-generated image crops from the bounding boxes to generate per-object masks. Contact-GraspNet then uses these masks to filter grasps that are not on objects of interest and keeps only a pruned set of grasps $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$. We visualize an example of this process in Fig. 2 for two different scenes. After pruning, we randomly select a single grasp pose, $\mathbf{T}_{\text{grasp}}$, from $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$. To approach without collisions, we include a pregrasp pose $\mathbf{T}_{\text{pregrasp}}$ that is defined with a translational offset above the grasp pose.
@@ III-A2 Grasping
Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory. An example of a grasp sequence is shown in Fig. 1(b). During grasping, two modules play a critical role in the robustness of PvP: compliant control for grasping (CCG) and tactile regrasping (TR).
@@
We use a Cartesian impedance controller for manipulation. This choice allows us to set the level of compliance of the manipulator at various stages as needed. Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values. We call this CCG. The high compliance of the robot arm allows the manipulator and gripper to comply with the contact constraints and minimize the applied forces against the environment. Larger contact-constrained objects generate unnecessary reaction forces when gripped without compliance. Large forces, in turn, trigger emergency stops and can damage the environment and the robot. Specific to PvP, large reaction forces also disrupt and move the object relative to the gripper, ultimately leading to inaccurate placements. On the other hand, when grasping with compliance, the environment contact constraints guide the manipulator towards a natural object placement pose.
@@
Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip significantly once grasped. In the real world, the accumulation of errors from various sources (e.g., camera calibration, controller, learned grasp planner, and noise from the RGB-D sensor and manipulator encoders) can result in grasps that are unstable (e.g., shallow grasps) and not ideal for PvP. To mitigate this issue, we use tactile sensing—specifically, a visuotactile Finger-STS [37] sensor—to preemptively detect if a grasp is stable and perform a regrasp if needed. The Finger-STS is capable of providing multimodal feedback (i.e., both tactile and visual observations). To this end, we design our TR module to detect the contact surface area and calculate the relative change in end effector (EE) pose required to recover a larger contact surface and, thereby, a more stable grasp. We detect regions of contact based on marker displacements and the RGB image from the Finger-STS. If partial contact is detected, we calculate the difference between the ideal and detected contact region and use this to calculate the commanded change in EE pose. We note that a simpler and more general random regrasp strategy could also be used. An example of this is shown in Fig. 3. We visualize contact regions for various household items in Fig. 4.
@@
In Section IV-A, we experimentally investigate the effects of including these modules through an ablation study.
III-A3 Retrieving
@@
We begin the retrieval phase when a valid, stable grasp is detected. An image of the retrieval procedure is shown in Fig. 1(c). We store the poses of the EE as well as the respective timestamps throughout the retrieval process. For our work, we measure joint encoder readings at a rate of 120 $\mathrm{H}\mathrm{z}$. The manipulator is set to be compliant along the rotational axes of the EE only, when commanded to move to the pregrasp pose. Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{\text{clearance}}$ around a set fixed pose above the scene. That is, given a set fixed pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$ to recover a clearance pose. We use $\sigma_{tr}$ = 2.5 cm during data collection.
@@
At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose,
@@
where $\mathbf{T}_{0}=\mathbf{T}_{\text{grasp}}$, $\mathbf{T}_{M}=\mathbf{T}_{\text{clearance}}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0. To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a sparser trajectory $\bar{\tau}^{\text{expert}}_{r}$. We extract a total of $\bar{M}+1$ states, where $\bar{M}=\lceil\frac{t_{M}-t_{0}}{\Delta t}\rceil$, by finding the nearest pose (in terms of time) from the dense trajectory $\tau^{\text{expert}}_{r}$ at every $\Delta t$ interval:
@@
In this work, our visual policies operate at 5 $\mathrm{H}\mathrm{z}$, so we set $\Delta t=$ 0.20 s.
III-A4 Placing by Reversing
We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory,